five years ago a woman age was the sqouare of her son age.ten years hence the mother age was twice of his son age 1 find the present age of woman 2 the age of son five years ago

Answer by ankor@dixie-net.com(22740)

five years ago a woman age was the square of her son age.
ten years hence the mother age was twice of his son age.
:
Let m = mother's age
Let s = son's age
:
m - 5 = (s-5)^2
m - 5 = s^2 - 10s + 25
m = s^2 - 10s + 25 + 5
m = s^2 - 10s + 30
:
We also know,"ten years hence the mother age was twice of his son age"
m + 10 = 2(s+10)
m + 10 = 2s + 20
m = 2s + 20 - 10
m = 2s + 10
therefore
s^2 - 10s + 30 = 2s + 10
s^2 - 10s - 2s + 30 - 10 = 0
A quadratic equation
s^2 - 12s + 20 = 0; Corrected a mistake here, Sorry!
Factors to
(s-10)(s-2) = 0
The only solution that will make sense here.
s = 10 yrs old is son's age
:
I'll let you find Mom's age, check solutions in both statements.